# Estimating the Number of HTTP/3 Responses in QUIC Using Deep Learning

### v2

@@ V-A Results for Known Web Servers
We first evaluate our framework when the web servers are known. Each server’s traces were split $80:20$ into training and testing sets, and we trained five models on different random splits. During training we used a batch size of 64, the Adam optimizer [36], and a ReduceLROnPlateau scheduler that reduced the learning rate by 30% upon reaching a validation-loss plateau. Early stopping with a patience of six epochs prevented overfitting. We performed a grid search to determine the optimal loss parameters $\alpha,\beta,\gamma$ by selecting the combination that yielded the lowest validation loss. We tested $\alpha\in\{0,0.3,0.5,0.7,1\}$, $\beta\in\{0,0.4,0.6,1\}$, and $\gamma\in\{1,2,3\}$. For $T=0.1$ second windows, $\alpha=0.7$, $\beta=0.4$, and $\gamma=2$ were optimal; for $T=0.1$ second windows, $\gamma=3$ worked best with the same $\alpha$ and $\beta$.
@@ V-B Results for Unknown Web Servers
Figure 6: Unknown-server prediction errors ($T=0.3$) with independent.co.uk” and google.com” as test servers. Red: median; blue boxes: 25–75% intervals.
@@ VI Evaluating on Complete Traces
Figures 7(7(a)) and 7(7(b)) illustrate a notable difference in predictive behavior between models that were trained and evaluated with $T=0.3-$ and $T=0.1-$second window sizes. Specifically, they show the presence of diagonal patterns in the predictions of the $T=0.3$ model on the test set that are absent in the $T=0.1$ predictions. This phenomenon exists for several reasons: (1) When using a $T=0.1$ subdivision, a very high percentage of the images’ true labels have lower class values, and the model that was trained using the $T=0.1$ window dataset is very accurate for low value classes, whereas a $T=0.3$ subdivision yields images with higher class values, hence increasing the variance of the true labels, when both models perform worse for the higher value classes as opposed to the lower class value; and (2) any incorrect prediction by either model contributes to an increase in the cumulative predictions for the remainder of the considered trace, thereby elevating the overall predicted values.
Both figures show that choosing the optimal window length is difficult and use-case-dependent. While shorter windows improve overall accuracy in cases where an aggregation of images is more important and offline analysis is available, longer windows improve the accuracy when examined on a single image (as seen in Section V). Using shorter windows, however, exacts a higher training and inference cost.

### v3

@@ I Introduction
While counting HTTP/3 responses might appear simpler than comprehensive QUIC traffic classification, traditional non-ML methods can struggle with parallel HTTP/3 streams and overlapping packet patterns in a single QUIC connection. The ML-based approach in DecQUIC exploits fine-grained, directional, and density features in a scalable way, improving on hand-crafted heuristics by learning subtle temporal and length-based signatures. Moreover, as QUIC evolves and server behaviors diversify, ML models can adapt more flexibly than fixed rules, ensuring our method remains robust across a wide range of deployment scenarios.
@@ II Related Work
In contrast to these existing classification efforts, our work addresses the subtler challenge of counting concurrent HTTP/3 responses in scenarios with highly imbalanced classes and partial concurrency. Most prior methods assume a limited server set (often Google-centric) and do not explicitly tackle out-of-distribution servers or the fine-grained concurrency within modern QUIC streams. By focusing on response counts, we enable tasks such as load balancing and attack detection, where concurrency plays a crucial role.
@@ V-A Results for Known Web Servers
We first evaluate our framework when the web servers are known. Each server’s traces were split $80:20$ into training and testing sets, and we trained five models on different random splits. During training we used a batch size of 64, the Adam optimizer, and a ReduceLROnPlateau scheduler that reduced the learning rate by 30% upon reaching a validation-loss plateau. Early stopping with a patience of six epochs prevented overfitting. We performed a grid search to determine the optimal loss parameters $\alpha,\beta,\gamma$ by selecting the combination that yielded the lowest validation loss. We tested $\alpha\in\{0,0.3,0.5,0.7,1\}$, $\beta\in\{0,0.4,0.6,1\}$, and $\gamma\in\{1,2,3\}$. For $T=0.1$ second windows, $\alpha=0.7$, $\beta=0.4$, and $\gamma=2$ were optimal; for $T=0.1$ second windows, $\gamma=3$ worked best with the same $\alpha$ and $\beta$.
@@ V-B Results for Unknown Web Servers
Figure 6: Unknown-server prediction errors ($T=0.1$) with independent.co.uk” and google.com” as test servers. Red: median; blue boxes: 25–75% intervals.
@@ VI Evaluating on Complete Traces
Figures 7(7(a)) and 7(7(b)) illustrate a notable difference in predictive behavior between the models trained and evaluated with $T=0.1$- and $T=0.3$-second window sizes. In particular, the $T=0.3$ model displays a series of strong diagonal patterns, whereas these are far less evident for the $T=0.1$ model. Several factors drive this phenomenon:
  1. 1.
    Variance in true labels: When using $T=0.1$-second windows, the majority of images fall into lower-valued classes (e.g., $0\!-\!3$ responses). Consequently, the model trained on such a distribution excels in this narrow lower-range. By contrast, a $T=0.3$-second window lumps together more concurrent requests, producing a broader range of labeled classes (including mid-to-high values). As a result, any systematic errors the model makes on higher counts tend to repeat across many samples, appearing as clustered diagonals in the scatter plot.
  2. 2.
    Cumulative effect of mispredictions: When summing per-window predictions to obtain the total number of responses per trace, even minor window-level misestimates can compound. If a single window is misestimated by a few counts (especially in mid-to-high classes), subsequent partial errors accumulate, shifting the aggregated prediction upward or downward. This systematic drift manifests visually as diagonal streaks above or below the perfect-prediction line.
  3. 3.
    Class imbalance at high values: Although $T=0.3$-second windows increase the likelihood of higher-class labels, those classes are still relatively rare compared to classes near the mean. The model’s exposure to fewer high-class examples during training can lead to amplified inaccuracies on that subset, reinforcing the diagonal banding once these errors propagate across entire traces.
  4. 4.
    Coarser temporal granularity: Finally, with larger windows, nuanced changes in packet arrival rates are “averaged out” in each image, potentially masking fine-grained cues that help distinguish, say, 10 from 12 concurrent responses. In contrast, $T=0.1$-second windows offer more granular snapshots, which, while increasing computational cost, can bolster per-trace accuracy by limiting each misestimation’s scope.
Both figures highlight that the choice of window length is inherently use-case-dependent. Shorter windows typically enhance aggregated trace-level accuracy—particularly for offline analyses that require precise counts—whereas longer windows can yield better single-image accuracy (as seen in Section V). However, the finer temporal resolution of $T=0.1$ seconds also increases computational overhead during both training and inference. Ultimately, balancing these trade-offs is essential for practitioners deploying DecQUIC in diverse QUIC environments.